OpenStudy (anonymous):

What is the value of the discriminant of 3x^2+3x+6=0? A) -63 B) 54 C) 63 D) 81

OpenStudy (vishweshshrimali5):

Recall that determinant for quadratic equation, \(ax^2 + bx + c = 0\) is given by: \[D = \cfrac{-b \pm \sqrt{b^2 - 4ac}}{2a}\]

OpenStudy (vishweshshrimali5):

Oops sorry! It is: \[D = b^2 - 4ac\]

OpenStudy (vishweshshrimali5):

Now, compare \(3x^2 + 3x + 6 = 0\) with \(ax^2 + bx + c = 0\) What are the values of a,b and c?

OpenStudy (anonymous):

A is 3 and so is B, C is 6

OpenStudy (vishweshshrimali5):

Great!! Now, can you find the determinant ?

OpenStudy (anonymous):

So it would be (3)^2-4(3)(6)=0

OpenStudy (anonymous):

which is 9-72

OpenStudy (vishweshshrimali5):

Great! what should be the final answer then ?

OpenStudy (anonymous):

so its -63?

OpenStudy (vishweshshrimali5):

Yep!!!!!!

OpenStudy (anonymous):

awesome!!!! thx

OpenStudy (vishweshshrimali5):

your welcome :)